# Markovian Quantum Computation

### 04 Mar 2024 v1

@@ Abstract
We propose a general prototype of quantum computation based on the finding of many-body ground states. The scheme is inspired by cavity cooling, involving the emulation of a zero-temperature reservoir. Repeated discarding of ancilla reservoir qubits extracts the entropy of the system, driving the system towards its ground state. At the same time, measurement of the discarded qubits hints at the energy level structure of many-body systems. We show that quantum computation based on this Markovian process is equivalent in its computational power to the one based on quantum circuits. We exemplify the scheme with several illustrative use cases for combinatorial optimization problems, discussing the efficiency and the issue of local energy minima. We also discuss its application to the preparation of quantum many-body ground states for gapped systems.
@@ I Introduction
It is widely believed that quantum computers are not universally efficient in solving general ground-state problems, even for classical Hamiltonians. Nonetheless, we still need to find out specific classes of such problems that benefit from quantum computation and are yet practically useful. Also, there’s a need to devise diverse quantum algorithms for them, each having its unique advantages. We have a number of tools available for such purposes [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] (Ref. [21] contains a nice summary of earlier schemes). The most prominent is adiabatic quantum computation (AQC) [10, 11, 12]. The AQC exploits the adiabatic evolution of the instantaneous ground state when the Hamiltonian is adiabatically varied. The time scale of the operation is thus governed by the minimal spectral gap during the evolution. However, analyzing the spectral gap is extremely difficult, leaving the performance of AQC largely unknown. There are also hybrid approaches, such as quantum approximate optimization algorithms (QAOA) [14, 15, 13] and variational quantum eigensolvers (VQE) [16, 17, 18], aided by classical optimization subroutines.
@@
As mentioned, the difficulty in analyzing AQC arises as the initial and final Hamiltonians intermingle. On the other hand, the properties of the final Hamiltonian alone can be largely unveiled, even though its ground state remains unknown. Given this, it is natural to consider schemes to directly cool the final Hamiltonian system to its ground state. While earlier schemes based on dissipative engineering [19, 20, 21] appear to achieve this goal, a closer inspection reveals that they actually lack desirable properties as an ideal cooling method. To clarify this, it is instructive to discuss classical and quantum Hamiltonians separately. By the former, we mean those Hamiltonians diagonalized in the computational basis.
@@
For classical Hamiltonians, the earlier dissipative schemes do not provide any quantum advantage. There are a few fundamental reasons. They implement a completely-positive trace-preserving (CPTP) map that transforms the state $\rho$ to $\sum_{k}E_{k}\rho E_{k}^{\dagger}$ [22]. Here, each Kraus operator $E_{k}$ directly accesses the associated local part of the system, driving it into a certain direction. However, as the energy of the whole system is a global quantity, determining the optimal direction based solely on a small part is not possible (note that this was the very origin of the geometric frustration, rendering the problem intractable). For this reason, the state is randomized when the measurement projects the state into an undesired subspace. Note, however, that this entire process is essentially equivalent to a mere random guess. Furthermore, for classical Hamiltonians, they do not involve any operations in non-computational bases. This means that they are, in fact, identical to classical probabilistic models. In the case of quantum Hamiltonians, the problem is more involving: one need to find an appropriate basis containing the ground state. This is the place where the above CPTP map comes into play. Apart from the role of fixing the basis, however, the above mentioned limitations persist.
@@
An optimal cooling should involve quantum transitions that take place exclusively in the direction of lowering energy. Here, a mechanism to prevent heating can be ensured only by an external agent, which globally accesses the system, gathers its energy, and relaxes it into a zero-temperature Markovian reservoir. This process is exactly given by cavity cooling methods [23]. The aim of this paper is to materialize this concept as a general quantum computation model and discuss its efficiency in some illustrative scenarios. Our model can be paralleled with AQC for their similarity in motivation. On the other hand, unlike AQC that is a unitary process, our model exploits a Markovian process. In order to stress this and for the sake of brevity, we will refer our model to Markovian quantum computation (MQC) hereafter.
@@
Being a Markovian process, MQC inherits the intrinsic advantages of the earlier dissipative schemes, alleviating the demand of precise system controls to some extent. On top of this, MQC takes true quantum advantages. In fact, we show that MQC is equivalent to quantum circuit models in terms of computational power. This argument is similar to the established one regarding the computational equivalence between AQC and quantum circuit models [11]. All three frameworks are thus computationally equivalent. However, this assertion needs to be interpreted with care. For instance, the equivalence of AQC to MQC is established by translating the unitary evolution of AQC into a quantum circuit, which is then converted into a Hamiltonian for MQC. However, this resulting Hamiltonian differs from the original problem Hamiltonian of the AQC. It is unclear if the feasibility of ground-state preparation by AQC implies that of the same Hamiltonian by MQC, and vice versa.
@@
Any ground-state preparation scheme, including AQC, inevitably suffers from the issue of local energy minima [6, 24, 25]. In MQC, one can design tunneling transitions out of local minima if the information is given on their nature. This flexibility also constitutes the advantage of MQC. In practice, one could try various transition terms in the Hamiltonian, while observing the occurrence of the transition from the detection of cavity photons. This, in turn, provides a method to inspect the energy-level structure of the Hamiltonian.
@@
In this work, our main focus will be placed on MQC for combinatorial optimization problems. Such problems involve classical Hamiltonians and have a number of favorable features for MQC. We consider two extreme cases. The first case is free from local energy minima, but the transition rate is superpolynomially small in the system size. This case turns out to be equivalent to Grover’s quantum search algorithm. The second case is an opposite limit, wherein the transition rate is maximized, but local energy minima begins to pose challenges. In this case, high-order transitions, required to get out of local energy minima, become the bottleneck. These transitions require a time that is exponentially large in the order, which can increase with the system size in general. Finally, we briefly discuss the case of quantum Hamiltonians. This case is heavily system-dependent and harder to analyze. We argue that the spectral gap above the ground state determines the basic time scale of the cooling.
@@ II Notation and working principles
where $E(z)\leq E(z^{\prime})$ for $z<z^{\prime}$. Let us call $H_{P}$ a problem Hamiltonian. Our aim is to find the ground state $|0\rangle_{P}$. Transitions among the energy levels $|z\rangle_{P}$ are allowed by introducing in the Hamiltonian what we call a transition term $H_{T}$, which does not commute with $H_{P}$. $H_{T}$ should be designed to suit the need of the given problem. We will consider two examples below. Here, we assume the energy level spacings of $H_{P}$ are characterized by an energy scale $\Delta>0$, and so is the norm $\|H_{T}\|$. Our aim is to introduce a parameter $0<\lambda\ll\Delta$ to treat $\lambda H_{T}$ as a perturbation to $H_{P}$.
If we consider the transition problem described by Hamiltonian $H_{P}+\lambda H_{T}$, the transition from $|z\rangle$ to $|z^{\prime}\rangle$ is allowed only when $|E(z)-E(z^{\prime})|\lesssim|\langle z^{\prime}|\lambda H_{T}|z\rangle|$, ignoring the influence of the degeneracy. In order to ensure $E(z^{\prime})<E(z)$, we introduce cavity field $c_{m}$ with resonant frequency $\omega_{m}\gg|\langle z^{\prime}|\lambda H_{T}|z\rangle|$ and couple it to the transition. Let us denote by $|\cdot\rangle_{m}$ the photon number state of the $m$-th cavity. If the cavity is initially empty, the transition from $|z\rangle_{P}|0\rangle_{m}$ to $|z^{\prime}\rangle_{P}|1\rangle_{m}$ is allowed only when $E(z^{\prime})\simeq E(z)-\omega_{m}$, providing a mechanism to avoid heating.
@@
Here, $a_{0}\in\{0,1\}$ determines whether the transitions preserving the cavity photon numbers are allowed. This term helps the state escaping from local minima through high-order transitions. In practice, it is generally enough to restrict the maximum number of photons in each cavity to one. In this case, $M$ cavity modes can be simulated with $M$ qubits.
Initially, the system is prepared in state $2^{-N/2}\sum_{z=0}^{2^{N}-1}|z\rangle_{P}\bigotimes_{m=1}^{M}|0\rangle_{m}$. A single cycle of cooling is performed by unitarily evolving the system according to Hamiltonian (2) for an appropriate duration of time, measuring the cavity states, and resetting every cavity to the vacuum state. Here, the measurement of the cavity photons is not mandatory, but it would reveal some information on the current state and the internal transition structure of the problem Hamiltonian.
The computation is done by repeating this cycle. The required number of cycles depends on the problem and strategy. One possible strategy is to repeat until the cavity photon is not detected for an enough number of cycles. This signals with high probability that the system has reached either the ground state or local minima. For NP problems that are guaranteed to have at least one solution, deciding which is the case is straightforward. Otherwise, one may repeat the same computation to increase the confidence. Given the unknown ground state not being verifiable in general, the decision is probabilistic after all.
@@ III Equivalence to circuit models
Figure 1: The transition diagram into which a quantum circuit is converted for MQC. $|\phi_{t}\rangle_{P}$, $|t\rangle_{C}$, and $|\cdot\rangle_{1}$ denotes, respectively, the state of the quantum circuit at the $t$-th time step, the clock state to label the time step, and the number state of the cavity. $\Delta$ is the energy difference between the adjacent levels and $\lambda$ is the Rabi frequency of the designated transition. The wavy lines represent the transition after detecting and clearing the cavity photon.
Ref. [11] shows that AQC is equivalent to the quantum circuit model in computational power up to polynomial overhead. The forward direction of the equivalence is rather straightforward because the entire procedure of AQC is a unitary transformation, which can be simulated efficiently in quantum circuits. Showing the other direction replies on the clock state $|t\rangle_{C}$ that is tagged to the state at the $t$-th time step of the quantum-circuit computation.
A similar equivalence relation can be shown for MQC. Again, showing that MQC can be simulated efficiently in quantum circuits is straightforward. We thus focus on the other direction. Suppose a certain quantum algorithm is run on a quantum circuit in $T$ time steps, where $U_{t}^{P}$ is applied to the qubits at the $t$-th time step. Let $|\phi_{t}\rangle_{P}\equiv U_{t}^{P}U_{t-1}^{P}\cdots U_{1}^{P}|\phi_{0}\rangle_{P}$ be the state after applying $U_{t}^{P}$, and tag this state with the clock state as $|\phi_{t}\rangle_{P}|t\rangle_{C}$. The idea is to get $|\phi_{t}\rangle_{P}|t\rangle_{C}$ transformed into $|\phi_{t+1}\rangle_{P}|t+1\rangle_{C}$ in the course of an energy-lowering transition. This is enabled by choosing the terms in Hamiltonian (2) as
@@
IV Combinatorial optimization by Markovian quantum computation
@@ IV.1 Notation
Consider combinatorial optimization problems that aim to find out the optimal solutions from a set of $2^{N}$ possible configurations represented by $N$-bit string $z\equiv z_{1}z_{2}\cdots z_{N}$. The optimality of each configuration is evaluated by a non-negative cost function $E(z)$. The objective is to identify configurations $z$ that minimize $E(z)$. This problem is straightforwardly translated into the ground-state finding problem, where the cost function $E(z)$ is identified with the energy eigenvalue in Hamiltonian (1). Here, we set $E(z)$ to be integer multiples of a parameter $\Delta>0$. The cavity frequencies in Hamiltonian (2) can then be chosen as integer multiples of $\Delta$. By choosing $M$ to be the maximum of $E(z)$, all possible transitions in the problem Hamiltonian can be coupled to a cavity transition. This imposes a polynomial overhead to the computational resource.
@@ IV.2 Reproduction of Grover’s search algorithm
To perform an analogous task in MQC, we adjust the Hamiltonian (2) as
@@
In this basis, the off-diagonal element becomes ${}_{P}\langle\phi_{0}|\lambda\bigotimes_{i}\frac{I_{i}+X_{i}}{2}|\phi_{1}\rangle_{P}\simeq\lambda n_{0}^{1/2}2^{-N/2}$. The computation time is thus $\mathcal{O}(\sqrt{2^{N}/n_{0}})$, identical to that of Grover’s algorithm. One advantage of MQC implementation is that the end of computation is signaled by a detection of the cavity photon. This is especially beneficial when the number of solutions $n_{0}$ is unknown.
In this example, the problem Hamiltonian has only two energy levels corresponding to the solution states and the rest. A naturally following question is if the performance can be improved when the problem Hamiltonian has more than two energy levels. It can be seen that with the transition term chosen above, more energy levels just make the computation slower. To see this, note that the quadratic speedup comes from the collective transition, the rate of which increases with the degeneracy of the energy level. The increase in the number of energy levels leads to the decrease in the degeneracy of each level, which diminishes the collective effect. Moreover, it results in an increased number of transitions required to reach the ground state.

### 08 Mar 2024 v2

# Quantum Computation by Cooling

@@ Abstract
Adiabatic quantum computation is a paradigmatic model aiming to solve a computational problem by finding the many-body ground state encapsulating the solution. However, its use of an adiabatic evolution depending on the spectral gap of an intricate many-body Hamiltonian makes its analysis daunting. While it is plausible to directly cool the final gapped system of the adiabatic evolution instead, an analysis of such a scheme is surprisingly missing. Here, we propose a specific Hamiltonian model for this purpose. The scheme is inspired by cavity cooling, involving the emulation of a zero-temperature reservoir. Repeated discarding of ancilla reservoir qubits extracts the entropy of the system, driving the system toward its ground state. At the same time, the measurement of the discarded qubits hints at the energy level structure of the system as a return. We show that quantum computation based on this cooling procedure is equivalent in its computational power to the one based on quantum circuits. We then exemplify the scheme with a few illustrative use cases for combinatorial optimization problems. In the first example, the cooling is free from any local energy minima, reducing the scheme to Grover’s search algorithm with a few improvements. In the second example, the cooling suffers from abundant local energy minima. To circumvent this, we implant a mechanism in the Hamiltonian so that the population trapped in the local minima can tunnel out by high-order transitions. We support this idea with a numerical simulation for a particular combinatorial optimization problem. We also discuss its application to preparing quantum many-body ground states, arguing that the spectral gap is a crucial factor in determining the time scale of the cooling.
@@ I Introduction
Identifying the ground states of quantum many-body Hamiltonians and understanding their characteristics are fundamental in physics. The ground states encode the spectral properties of the Hamiltonian [1, 2, 3, 4, 5, 6, 7] and represent the essential quantumness of the system. The latter becomes particularly evident in gapped systems, in which the ground state predominantly determines the entire properties at sufficiently low temperature. This lays the foundation for various exotic quantum phenomena, such as quantum Hall effects and superconductivity [8].
@@
It is widely believed that quantum computers are not universally efficient in solving general ground-state problems, even for classical Hamiltonians. Nonetheless, we still need to find out specific classes of such problems that benefit from quantum computation and are yet practically useful. Also, there’s a need to devise diverse quantum algorithms for them, each having its unique advantages. We have a number of tools available for such purposes [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29] (Ref. [29] contains a nice summary of earlier schemes). The most prominent is adiabatic quantum computation (AQC) [18, 19, 20]. The AQC exploits the adiabatic evolution of the instantaneous ground state when the Hamiltonian is varied. The time scale of the operation is thus governed by the minimal spectral gap during the evolution. However, analyzing the spectral gap is extremely difficult, leaving the performance of AQC largely unknown. There are also hybrid approaches, such as quantum approximate optimization algorithms (QAOA) [22, 23, 21] and variational quantum eigensolvers (VQE) [24, 25, 26], aided by classical optimization subroutines.
@@
As mentioned, the difficulty in analyzing AQC arises as the initial and final Hamiltonians intermingle. On the other hand, the properties of the final Hamiltonian alone can be largely unveiled, even though its ground state remains unknown. Given this, it is natural to consider schemes to directly cool the final Hamiltonian system to its ground state. While earlier schemes based on dissipative engineering [27, 28, 29] seem to achieve this goal, a closer inspection reveals that they actually lack desirable properties as a cooling method. To clarify this, it is instructive to discuss classical and quantum Hamiltonians separately. By the former, we mean those Hamiltonians diagonalized in the computational basis.
@@
For classical Hamiltonians, the earlier dissipative schemes are reduced to a classical random search. There are a few fundamental reasons. They implement a completely-positive trace-preserving (CPTP) map that transforms the state $\rho$ to $\sum_{k}E_{k}\rho E_{k}^{\dagger}$ [30]. Here, each Kraus operator $E_{k}$ directly accesses the associated local part of the system, driving it into a certain direction. However, as the energy of the whole system is a global quantity, determining the optimal direction based solely on a small part is not possible (note that this was the very origin of the geometric frustration, rendering the problem intractable). For this reason, the state is randomized when the measurement projects the state into an undesired subspace. Note, however, that this entire process is essentially equivalent to a mere random guess. Furthermore, for classical Hamiltonians, they do not involve any operations in non-computational bases. This means that they are, in fact, identical to classical probabilistic models. In the case of quantum Hamiltonians, the problem is more involving: one needs to find an appropriate basis containing the ground state. This is the place where the above CPTP map comes into play. Apart from the role of fixing the basis, however, the above-mentioned limitations persist.
@@
Simply put, both cooling and heating coexist in the above-mentioned schemes. A genuine cooling scheme should rely on quantum transitions that take place exclusively in the direction of lowering energy. It is important to understand that a heating-prevention mechanism requires an external agent, which globally accesses the system, gathers its energy, and relaxes it into a Markovian reservoir. This process is exactly provided by cavity cooling methods [31]. This paper aims at materializing this concept as a general quantum computation model and discuss its efficiency in some illustrative scenarios. Our model can be paralleled with AQC for their similarity in motivation. On the other hand, unlike AQC, our model exploits a non-unitary process. For brevity, we will refer to our model as cooling-based quantum computation (CQC) hereafter.
@@
Being a dissipative process, CQC inherits the intrinsic advantages of the earlier dissipative schemes, alleviating the demand of precise system controls to some extent. On top of this, CQC takes true quantum advantages. In fact, we show that CQC is equivalent to quantum circuit models in terms of computational power. This argument is similar to the established one regarding the computational equivalence between AQC and quantum circuit models [19]. All three frameworks are thus computationally equivalent. However, this assertion needs to be interpreted with care. For instance, the equivalence of AQC to CQC is established by translating the unitary evolution of AQC into a quantum circuit, which is then converted into a Hamiltonian for CQC. However, this resulting Hamiltonian differs from the original problem Hamiltonian of the AQC. It is an open question whether the feasibility of ground-state preparation by AQC implies that of the same Hamiltonian by CQC, and vice versa.
@@
Any ground-state preparation scheme, including AQC, inevitably suffers from the issue of local energy minima [14, 32, 33]. In CQC, one can design tunneling transitions out of local minima if the information is given on their nature. This flexibility also constitutes the advantage of CQC. In practice, one could try various transition terms in the Hamiltonian, while observing the occurrence of the transition by monitoring cavity photons. This, in turn, provides a method to inspect the energy-level structure of the Hamiltonian. For example, the presence or absence of a finite spectral gap above the ground state could be inspected for unsolvable Hamiltonians.
@@
In this work, much of our focus will be placed on CQC for combinatorial optimization problems, which have favorable features for CQC and allow for the analytical treatment. We consider two extreme cases. The first case is free from local energy minima, but the transition rate is superpolynomially small in the system size. This case turns out to be equivalent to Grover’s quantum search algorithm, aside from some additional advantages that CQC brings. The second case is an opposite limit, wherein the transition rate is maximized, but local energy minima begin to pose challenges. In this case, high-order transitions, which are required to get out of local energy minima, become the bottleneck. These transitions require a time exponentially large in the order, which can increase with the system size in general. We perform the numerical simulation for a particular combinatorial optimization problem to support this idea. Finally, we briefly discuss the case of quantum Hamiltonians. This case is heavily system-dependent and more complicated to analyze. We argue that the spectral gap above the ground state is a crucial factor in determining the time scale of the cooling. This is a natural consequence of the energy-time uncertainty principle.
@@ II Notation and working principles
where $E(z)\leq E(z^{\prime})$ for $z<z^{\prime}$. Let us call $H_{P}$ a problem Hamiltonian. We aim to find the ground state $|0\rangle_{P}$. Transitions among the energy levels $|z\rangle_{P}$ are allowed by introducing in the Hamiltonian what we call a transition term $H_{T}$, which does not commute with $H_{P}$. $H_{T}$ should be designed to suit the needs of the given problem. We will consider two simple examples below. Here, we assume the energy level spacings of $H_{P}$ are characterized by an energy scale $\Delta>0$, and so is the norm $\|H_{T}\|$. Our aim is to introduce a parameter $0<\lambda\ll\Delta$ to treat $\lambda H_{T}$ as a perturbation to $H_{P}$.
If we consider the transition problem described by Hamiltonian $H_{P}+\lambda H_{T}$, the transition from $|z\rangle$ to $|z^{\prime}\rangle$ is allowed only when $|E(z)-E(z^{\prime})|\lesssim|\langle z^{\prime}|\lambda H_{T}|z\rangle|$, where the influence of the degeneracy is ignored. In order to ensure $E(z^{\prime})<E(z)$, we introduce cavity field $c_{m}$ with resonant frequency $\omega_{m}\gg|\langle z^{\prime}|\lambda H_{T}|z\rangle|$ and couple it to the transition. Here, the subscript $m$ is the index for different cavities. Let us denote by $|\cdot\rangle_{m}$ the photon number state of the cavity. If the cavity is initially empty, the transition from $|z\rangle_{P}|0\rangle_{m}$ to $|z^{\prime}\rangle_{P}|1\rangle_{m}$ is allowed only when $E(z^{\prime})\simeq E(z)-\omega_{m}$, providing a mechanism to prevent heating.
@@
Here, $a_{0}\in\{0,1\}$ determines whether the transitions preserving the cavity photon numbers are allowed. This term helps the state escape from local minima through high-order transitions. In practice, it is generally enough to restrict the maximum number of photons in each cavity to one. In this case, $M$ cavity modes can be simulated with $M$ qubits.
Initially, the system is prepared in state $2^{-N/2}\sum_{z=0}^{2^{N}-1}|z\rangle_{P}\bigotimes_{m=1}^{M}|0\rangle_{m}$. A single cycle of cooling is performed by unitarily evolving the system according to Hamiltonian (2) for an appropriate duration of time, measuring the cavity states, and resetting every cavity to the vacuum state. Here, the measurement of the cavity photon is not mandatory, but it would reveal some information on the current state and the internal transition structure of the problem Hamiltonian.
The computation is done by repeating this cycle. The required number of cycles depends on the problem and strategy. One possible strategy is to repeat until the cavity photon is not detected for an enough number of cycles. This signals with high probability that the system has reached either the ground state or local minima. For NP problems that are guaranteed to have at least one solution, deciding which is the case is straightforward. Otherwise, one may repeat the same computation to increase the confidence. Given that the unknown ground state is not verifiable in general, the decision is probabilistic after all.
We will implicitly assume that the evolution by the Hamiltonian (2) is simulated on a universal quantum computer, which is reasonable given the complexity of the Hamiltonian. In our analysis below, the polynomial overhead associated with this implementation is not taken into account.
@@ III Equivalence to circuit models
Figure 1: The transition diagram into which a quantum circuit is converted for CQC. $|\phi_{t}\rangle_{P}$, $|t\rangle_{C}$, and $|\cdot\rangle_{1}$ denotes, respectively, the state of the quantum circuit at the $t$-th time step, the clock state to label the time step, and the number state of the cavity. $\Delta$ is the energy difference between the adjacent levels and $\lambda$ is the Rabi frequency of the designated transition. The wavy lines represent the transition after detecting and clearing the cavity photon.
Ref. [19] shows that AQC is equivalent to the quantum circuit model in computational power up to polynomial overhead. The forward direction of the equivalence is straightforward because the entire procedure of AQC is a unitary transformation, which can be simulated efficiently in quantum circuits. Showing the other direction relies on the clock state $|t\rangle_{C}$ tagged to the state at the $t$-th time step of the quantum-circuit computation.
A similar equivalence relation can be shown for CQC. Again, showing that CQC can be simulated efficiently in a quantum circuit is straightforward. We thus focus on the other direction. Suppose a quantum algorithm is run on a quantum circuit in $T$ time steps, where $U_{t}^{P}$ is applied to the qubits at the $t$-th time step. Let $|\phi_{t}\rangle_{P}\equiv U_{t}^{P}U_{t-1}^{P}\cdots U_{1}^{P}|\phi_{0}\rangle_{P}$ be the state right after applying $U_{t}^{P}$, and tag this state with the clock state as $|\phi_{t}\rangle_{P}|t\rangle_{C}$. The idea is to get $|\phi_{t}\rangle_{P}|t\rangle_{C}$ transformed into $|\phi_{t+1}\rangle_{P}|t+1\rangle_{C}$ in the course of an energy-lowering transition. This is enabled by choosing the terms in Hamiltonian (2) as
@@
IV Combinatorial optimization by cooling-based quantum computation
@@ IV.1 Notation
Consider combinatorial optimization problems that aim to find the optimal solutions from a set of $2^{N}$ possible configurations represented by $N$-bit string $z\equiv z_{1}z_{2}\cdots z_{N}$. The optimality of each configuration is evaluated by a non-negative cost function $E(z)$. The objective is to identify configurations $z$ that minimize $E(z)$. This problem is straightforwardly translated into the ground-state finding problem, where the cost function $E(z)$ is identified with the energy eigenvalue in Hamiltonian (1). Here, we set $E(z)$ as integer multiples of a parameter $\Delta>0$. For example, for boolean satisfiability (SAT) problems, $E(z)$ is the number of unsatisfied clauses by the configuration $z$ [12]. The cavity frequencies in Hamiltonian (2) can then be chosen as integer multiples of $\Delta$. By choosing $M$ to be the maximum of $E(z)$, all possible transitions in the problem Hamiltonian can be coupled to a cavity transition.
@@ IV.2 Reproduction of Grover’s search algorithm
To perform an analogous task in CQC, we adjust the Hamiltonian (2) as
@@
In this basis, the off-diagonal element becomes ${}_{P}\langle\phi_{0}|\lambda\bigotimes_{i}\frac{I_{i}+X_{i}}{2}|\phi_{1}\rangle_{P}\simeq\lambda n_{0}^{1/2}2^{-N/2}$. The computation time is thus $\mathcal{O}(\sqrt{2^{N}/n_{0}})$, identical to that of Grover’s algorithm. One advantage of CQC implementation is that the end of computation is signaled by a detection of the cavity photon. This is especially beneficial when the number of solutions $n_{0}$ is unknown.
In this example, the problem Hamiltonian has only two energy levels corresponding to the solution states and the rest. A naturally following question is if the performance can be improved when the problem Hamiltonian has more than two energy levels. It can be seen that with the transition term chosen above, more energy levels just make the computation slower. To see this, note that the quadratic speedup comes from the collective transition, the rate of which increases with the degeneracy of the energy level. The increase in the number of energy levels leads to a decrease in the degeneracy of each level, which diminishes the collective effect. Moreover, it results in an increased number of transitions required to reach the ground state.